OpenStudy (anonymous):

what dy/dx for y=cot^2 (e^3x)

OpenStudy (anonymous):

Note: \(\frac{d}{dx} (cot x)=-csc x cot x\) \(\frac{d}{dx} (e^x)=e^x\) Are you familiar with the chain rule and power rule?

OpenStudy (anonymous):

yes i familia with that.. i get -3cosec^2 e^3x? what do you think?